# A Geometry-Grounded Data Perimeter in Azure

### v3

@@ 4.1 Formulation
As will be shown in the ultrametric TSP pseudo-code below, the Data Perimeter is not only geometrically grounded but also computationally tractable in our setting.
@@ 6 Conclusion
Given an ultrametric space $(X,d)$ with $n$ points, an ultracycle is a configuration where all pairwise distances are equal: $d(x_{i},x_{j})=\delta$ for all distinct $i,j\in\{1,2,\ldots,n\}$ and some constant $\delta>0$.
@@
Suppose all pairwise distances are equal: $d(x_{i},x_{j})=\delta$ for all $i\neq j$.
@@
$P_{\text{ultra}}=\sum_{i=1}^{n}d(x_{\pi(i)},x_{\pi(i+1)})=\sum_{i=1}^{n}\delta=n\delta$ (1)
where $x_{\pi(n+1)}:=x_{\pi(1)}$.
Since all edges have the same weight $\delta$, every possible tour has identical length $n\delta$.
@@
$d(x_{i},x_{j})\neq d(x_{k},x_{\ell})$ (2)
Let $d_{\min}=\min_{i\neq j}d(x_{i},x_{j})$ and $d_{\max}=\max_{i\neq j}d(x_{i},x_{j})$.
By assumption, $d_{\min}<d_{\max}$.
Lemma: in any ultrametric space, if $d(a,b)<d(a,c)$, then $d(b,c)=d(a,c)$.
Proof of Lemma: By the strong triangle inequality: $d(a,c)\leq\max\{d(a,b),d(b,c)\}$.
So we must have $d(a,c)\leq d(b,c)$.
Similarly, $d(b,c)\leq\max\{d(b,a),d(a,c)\}$. But $d(a,b)<d(a,c)$, so $d(b,c)\leq d(a,c)$.
Therefore, $d(b,c)=d(a,c)$. This proves the Lemma.
Now, let $S_{\min}=\{(i,j):d(x_{i},x_{j})=d_{\min}\}$ be the set of point pairs with minimum distance.
@@
Proof of Claim: Suppose, for contradiction, that there exists a tour using only edges of distance $d_{\min}$. This would mean we can connect all $n$ points using only edges from $S_{\min}$, forming a Hamiltonian cycle.
However, by the ultrametric property (Lemma), if $d(x_{i},x_{j})=d_{\min}$ and $d(x_{j},x_{k})=d_{\min}$, then either:
  • $d(x_{i},x_{k})=d_{\min}$ (all three points are equidistant), or
  • $d(x_{i},x_{k})>d_{\min}$ (the points form a non-equidistant triple)
If all points were connected by minimum-distance edges, then by transitivity of the equidistance relation in ultrametric spaces, all pairwise distances would equal $d_{\min}$, contradicting our assumption that the configuration is not an ultracycle.
Therefore, any Hamiltonian cycle must include at least one edge with distance $>d_{\min}$.
Completing the Proof:
Let $P_{\text{non-ultra}}$ be the length of an optimal TSP tour in the non-ultracycle configuration.
Since the tour must include at least one edge of length $>d_{\min}$, and the remaining $(n-1)$ edges have length $\geq d_{\min}$:
$P_{\text{non-ultra}}>d_{\min}+(n-1)\cdot d_{\min}=n\cdot d_{\min}$ (3)
For an ultracycle with the same minimum distance $d_{\min}$ (i.e., $\delta=d_{\min}$):
$P_{\text{ultra}}=n\cdot d_{\min}$ (4)
Therefore: $P_{\text{non-ultra}}>P_{\text{ultra}}$.
This proves that ultracycles uniquely minimize the TSP tour length in ultrametric spaces.

### v4

@@ 4.1 Formulation
As shown in the ultrametric TSP pseudo-code implementing the nearest-neighbor algorithm below, the Data Perimeter is not only geometrically grounded but also computationally tractable in our setting.
@@ 6 Conclusion
Given an ultrametric space $(X,d)$ with $n$ points, an ultracycle is a configuration where all pairwise distances are equal: $d(x_{i},x_{j})=\xi$ for all distinct $i,j\in\{1,2,\ldots,n\}$ and some constant $\xi>0$.
@@
Suppose all pairwise distances are equal: $d(x_{i},x_{j})=\xi$ for all $i\neq j$.
@@
$P_{\text{ultra}}=\sum_{i=1}^{n}d(x_{\pi(i)},x_{\pi(i+1)})=\sum_{i=1}^{n}\xi=n\xi,\text{ where }x_{\pi(n+1)}:=x_{\pi(1)}.$
Every possible tour has identical length $n\xi$.
@@
$d(x_{i},x_{j})\neq d(x_{k},x_{\ell})$
Let $d_{\min}=\min_{i\neq j}d(x_{i},x_{j})$.
@@
Proof of Claim:
Suppose, for contradiction, that there exists a tour using only edges of distance $d_{\min}$.
For any three points $x_{i}$, $x_{j}$, $x_{k}$:
  • If $d(x_{i},x_{j})=d_{min}$ and $d(x_{j},x_{k})=d_{min}$, then $d(x_{i},x_{k})\leq\max\{d(x_{i},x_{j}),d(x_{j},x_{k})\}=d_{min}$
  • Since $d_{min}$ is the global minimum distance, $d(x_{i},x_{k})\geq d_{min}$
  • Thus, $d(x_{i},x_{k})=d_{min}$
By induction, any two points connected by a path of $d_{min}$ edges must be at distance $d_{min}$.
Since the tour connects all $n$ points via $d_{min}$ edges, all pairwise distances must be $d_{min}$, forming and ultracycle.
This contradicts the non-ultracycle assumption.
Ultracycles uniquely minimize the TSP tour length in ultrametric spaces.